OpenStudy (anonymous):

5x3 = 3/5 and 1/6 of 2

OpenStudy (solomonzelman):

3/5 hours = 1/6 of the distance. multiply both sides times 5/3

OpenStudy (solomonzelman):

you will see that 3/5 will cancel with 5/3 on the left side, so have 1 hour. what do you get on the right side, though?

OpenStudy (solomonzelman):

1/6 times 5/3 =? (you tell me)

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

so in 1 hour, you travel 5/18 th of the distance.

OpenStudy (solomonzelman):

yw